OpenStudy (anonymous):

Please Help l(n)=1700X0.02n Find n if l(n)=136

OpenStudy (yttrium):

Just do the algebra 1st to find the value of your I(n) and then substitute I(n) to 136, then you will arrive at an answer.

OpenStudy (anonymous):

would I do l(n)=34n ? @Yttrium

OpenStudy (zzr0ck3r):

what is the X?

OpenStudy (yttrium):

Yes. :)

OpenStudy (zzr0ck3r):

that was a question

OpenStudy (zzr0ck3r):

\[l(n) = 1700*0.02n\]we want to know what is n when\[l(n) = 136\]so set l(n) equa l to that\[1700*0.02n=136\\n=\frac{136}{1700*0.02}\]

OpenStudy (yttrium):

@najaas Did you get it?

OpenStudy (anonymous):

is it 4?

OpenStudy (yttrium):

Great! You got it! :)

OpenStudy (anonymous):

okay thank you! @Yttrium

OpenStudy (anonymous):

thank you @zzr0ck3r